1. A 6.075 gram sample of an organic compound containing C, H and O is analyzed by combustion analysis and 10.12 grams of CO2 and 3.314 grams of H2O are produced.

In a separate experiment, the molar mass is found to be 132.1 g/mol. Determine the empirical formula and the molecular formula of the organic compound.
Answer #1

first calculate % of each element .

44 g CO2 contain 12 g O

10.12 g CO2 contain 10.12 x 12 / 44 = 2.76 g C

% C = (2.76 / 6.075) x 100

% C = 45.43

18 g H2O contain 2 g H

3.314 g H2O contain 3.314 x 2 / 18 = 0.368 g H

% H = (0.368 / 6.075) x 100

% H = 6.06 %

% O = 100 - 45.43 - 6.06 = 48.51

% element atomic ratio simplest ration C = 45.43 45.43/12 = 3.78 3.78/3.03 = 1.25 x 4 = 5.0 H = 6.06 6.06/10 = 6.06 6.06 /3.0